# Performance analysis for high-dimensional Bell-state quantum illumination

### v1

@@ Abstract
Quantum illumination (QI) is an entanglement-based protocol for improving lidar/radar detection of unresolved targets beyond what a classical lidar/radar of the same average transmitted energy can do. Originally proposed by Lloyd as a discrete-variable quantum lidar, it was soon shown that his proposal offered no quantum advantage over its best classical competitor. Continuous-variable, specifically Gaussian-state, QI has been shown to offer true quantum advantage, both in theory and in table-top experiments. Moreover, despite its considerable drawbacks, the microwave version of Gaussian-state QI continues to attract research attention. Recently, however, Pannu et al. (arXiv:2407.08005 [quant-ph]) have: (1) combined the entangled state from Lloyd’s QI with the channel models from Gaussian-state QI; (2) proposed a new positive operator-valued measurement for that composite setup; and (3) showed that, unlike Gaussian-state QI, their QI achieves the Nair-Gu lower bound on QI target-detection error probability at all noise brightnesses. Pannu et al.’s analysis was asymptotic, i.e., it presumed infinite-dimensional entanglement. This paper works out the finite-dimensional performance of Pannu et al.’s QI. It shows that there is a threshold value for the entangled-state dimensionality below which there is no quantum advantage, and above which the Nair-Gu bound is approached asymptotically. Moreover, in high-brightness noise, with both systems operating 1 dB below the Nair-Gu bound’s error-probability exponent, Pannu et al.’s QI requires much higher entangled-state dimensionality to achieve useful error probabilities than does Gaussian-state QI.
@@ I Introduction
Lloyd [5] coined the term “quantum illumination” for a lidar that transmitted a sequence of $M$-mode single-photon states while retaining their maximally-entangled single-photon companions for a joint measurement with the returned radiation. He assumed that the environment being probed never returned more than one photon in response to each transmission, and that photon was either background noise or a target reflection. Lloyd compared his QI performance with that of a single-photon (SP) lidar that probed the environment with the same state as his QI lidar, but had no stored idler. He argued that QI’s entanglement would make it difficult for a background photon to masquerade as the entangled companion of QI’s stored idler. Indeed, as compared to SP target detection, his QI target detection afforded a factor-of-$M$ improvement in error-probability exponent in the high-noise regime, i.e., when it is vastly more probable that the returned radiation from a single transmission is due to background as opposed to target reflection. That said, Lloyd’s QI and SP lidars are both quantum lidars, as they employ nonclassical transmitter states. Thus, when Shapiro and Lloyd [6] compared Lloyd’s QI with its best classical-lidar counterpart it turned out that the former could do no better than the latter, and could perform much worse. As a result, interest in Lloyd’s discrete-variable QI languished and was supplanted by interest in a continuous-variable version of QI, viz., Tan et al.’s Gaussian-state QI lidar [7].
@@
Pannu et al. do not assume that at most one photon is returned per transmission from the region being probed, hence avoiding the root cause of Lloyd’s QI not outperforming its best classical competitor. Their analysis is asymptotic, in that they pass to the limit $M\rightarrow\infty$ for the discrete-variable entangled state introduced in Lloyd’s QI. Doing so drives their receiver’s false-alarm probability to zero, and makes it simple to prove that their receiver realizes 6 dB quantum advantage in error-probability exponent in high-brightness noise, matching both Tan et al.’s performance in that regime and the Nair-Gu bound [13] on the attainable error-probability exponent of all possible QI protocols. Moreover, their receiver outperforms Gaussian-state QI for $\sim$1 photon/mode noise brightness, again matching the Nair-Gu bound for that operating regime. It is obvious that requiring infinite entangled-state dimensionality, i.e., $M\rightarrow\infty$, will put their system beyond the realm of practicality, and their paper presents no finite-$M$ results for the error-probability exponent.
This paper remedies the preceding problem. In particular, we introduce a new, more explicit, result for the joint density operator of the returned and retained radiation when the target is present. Using this result we obtain the single-shot false-alarm and detection probabilities for Pannu et al.’s receiver, and from those answers we get the finite-$M$ multi-shot likelihood-ratio test and use the Chernoff bound to get its error-probability exponent. We find that Pannu et al.’s QI has good and bad regimes, analogous to those of Lloyd’s QI, and that only in the good regime—which requires $M$ to exceed a threshold value—does it offer quantum advantage for the error-probability exponent.
@@
where the $\{\hat{a}_{B_{m}}:1\leq m\leq M\}$ are annihilation operators for background-noise modes that are in independent, identically-distributed (iid) thermal states with average photon number $N_{B}$. On the other hand, under hypothesis $H_{1}$ (target present), they use
@@
Pannu et al.’s QI thus offers a 6 dB quantum advantage in error-probability exponent when $N_{B}\gg 1$, no quantum advantage when $N_{B}\ll 1$, and 4.6 dB quantum advantage at $N_{B}=1$. These results match the $N_{B}\gg 1$ and $N_{B}\ll 1$ performance of Tan et al.’s QI, but that system only realizes a 2.8 dB quantum advantage at $N_{B}=1$ [14]. See Sec. VI for a more detailed appraisal of Pannu et al. QI versus Tan et al. QI.
@@ V Multi-Shot Likelihood Ratio and Error-Probability Exponent
$\exp$$\{-[N_{T}\kappa/(N_{B}+1)]\mathcal{E}(\kappa,N_{B},M)\}$
@@
and we will focus our attention on its $\kappa\ll 1$, $N_{B}\gg 1$, behavior, i.e., microwave operation, in which QI offers its maximum quantum advantage.
It is easily verified that $\lim_{M\rightarrow\infty}\mathcal{E}(\kappa,N_{B},M)=1$, so that our error-probability exponent matches the Nair-Gu bound at infinite $M$, as shown by Pannu et al. But how high must $M$ be to approach that limit? Figure 1 is a plot of $\mathcal{E}(\kappa,N_{B},M)$ versus $\log_{10}(M)$ for the representative case of $\kappa=0.001$ and $N_{B}=100,$ and Table 1 lists some key values therefrom. We see that: (1) below the $M=7.34\times 10^{5}$ threshold, Pannu et al. QI offers no quantum advantage in error-probability exponent when $\kappa=0.001$ and $N_{B}=100$; (b) $M=1.97\times 10^{7}$ is required for Pannu et al. QI to achieve 3 dB quantum advantage when $\kappa=0.001$ and $N_{B}=100$, i.e., the quantum advantage that Tan et al. QI realizes with its suboptimal parametric-amplifier receiver architecture [16]; and (3) $M=2.27\times 10^{13}$ is necessary for Pannu et al. QI to be 1 dB below the Nair-Gu error-probability exponent when $\kappa=0.001$ and $N_{B}=100$.
Figure 1: Plot of the penalty function $\mathcal{E}(\kappa,N_{B},M)$ versus $\log_{10}M$ for $\kappa=0.001$ and $N_{B}=100$. Note that the $M$ values shown here all exceed the $M_{0}=31.5$ threshold for $p_{D}>p_{F}$ when $\kappa=0.001$ and $N_{B}=100$.
@@
Table 1: $M$ values needed for representative values of $\mathcal{E}(0.001,100,M)$.
@@ VI Conclusions and Discussion
At this juncture, a comparison between finite-$M$ Pannu et al. QI and Tan et al. QI is warranted. Both systems can match the Nair-Gu bound on target-detection error-probability exponent for a weakly-reflecting ($\kappa\ll 1$) target embedded in high-brightness ($N_{B}\gg 1$) background noise. Moreover, neither offers any quantum advantage for low-brightness ($N_{B}\ll 1$) background noise. At $N_{B}=1$, Pannu et al. QI can offer a $4.6$ dB quantum advantage, whereas Tan et al. QI only affords $\sim$3 dB quantum advantage. That said, the conditions required for each of these protocols to realize their respective quantum advantages are quite different.
@@
Consider Pannu et al. QI and Tan et al. QI for our $\kappa=0.001$, $N_{B}=100$ example, with both systems operating at 1 dB below the Nair-Gu error-probability exponent. Pannu et al.’s QI requires $M=2.27\times 10^{13}$ to operate at 1 dB below the Nair-Gu error-probability exponent, and achieves the Chernoff-bound performance shown in Fig. 2. Tan et al.’s QI, on the other hand, requires the signal brightness to be $N_{S}=0.01523$ to operate at 1 dB below the Nair-Gu bound, and it achieves Fig. 2’s Chernoff-bound performance, where $N_{T}=MN_{S}$ is now the average number of transmitted signal photons. Figure 3 compares the entanglement dimensionalities of the two QI systems for the parameters used in Fig. 2. Here we see that Pannu et al. QI requires more than $10^{5}$ times the dimensionality that suffices for Tan et al. QI.
@@
Figure 2: Log-log plot of the Chernoff bound, $\Pr(e)$, versus $N_{T}$ for Pannu et al. QI and Tan et al. QI when $\kappa=0.001$, $N_{B}=100$ and both systems are operating at 1 dB below the Nair-Gu error-probability exponent. For Pannu et al. QI, this operating point requires $M=2.27\times 10^{13}$ and $N_{T}$ is the number of transmitted signal photons. For Tan et al. QI, this operating point requires $N_{S}=0.01523$ and $N_{T}$ is the average number of transmitted signal photons.
@@
Figure 3: Log-log plot of $M_{\rm Pannu}/M_{\rm Tan}$ versus $\Pr(e)$ where $M_{\rm Pannu}$ and $M_{\rm Tan}$ are the entangled-state dimensionalities associated with the $\Pr(e)$ versus $N_{T}$ curve from Fig. 2.
@@
As if Pannu et al. QI’s requiring $10^{5}$ times the entangled-state dimensionality of Tan et al. QI to achieve the same error probability were not bad enough, it incurs an even larger disadvantage when we look at time-bandwidth product. In the scalar-wave, unresolved-target scenario that we are considering, only temporal degrees of freedom are available. Hence, a single pulse of dimensionality $M$ must have a time duration $T$ and bandwidth $W$ satisfying $M=TW$. Reference [3] explains that Tan et al.’s QI can carve $T=M/W$-s duration pulses from the signal and idler outputs of a continuous-wave-pumped parametric downconverter with $W$-Hz phase-matching bandwidth to obtain the $M$-dimensional entangled state it needs. It is not clear how to generate the $M$-dimensional entangled-state pulses that Pannu et al. QI needs, but a sequence of $N_{T}>10^{5}$ such pulses are needed for that QI system to realize the Chernoff-bound error probabilities from Fig. 2. Thus Pannu et al. QI’s total required time-bandwidth product is at least $10^{10}$ times what suffices for Tan et al.’s QI in our Fig. 2 example.
@@
Our final points of comparison between these two QI protocols concern the four enormous hurdles, cited in Sec. I, that currently preclude finding a realistic target-detection use case for Tan et al.’s QI: its need to interrogate one resolution bin at a time; its need for a quantum memory to store its high time-bandwidth product idler; its need for extremely high time-bandwidth product radiation to get an acceptably low error probability; and its need for an interferometric measurement. How does Pannu et al.’s QI stack up against these hurdles? There is at least some encouragement in this regard, but so far the overall prospects are pretty poor. In particular, Pannu et al.’s QI still needs to interrogate one resolution bin at a time, and it still needs a quantum memory to store its high-dimensionality idler. However, assuming Pannu et al. QI’s single-pulse idler dimensionality is the same as that of Tan et al.’s QI, its discrete-variable memory may be easier to implement than Tan et al.’s continuous-variable memory. As we have seen, however, Pannu et al.’s required single-pulse idler dimensionality is apt to be $10^{5}$ times that of Tan et al.’s QI. Worse, Pannu et al. QI’s total time-bandwidth product may have to be a factor of $10^{10}$ times that of Tan et al. QI, owing to its need to transmit a long sequence of single-photon pulses. A final positive note for Pannu et al.’s QI is that it does not require an interferometric measurement. This phase insensitivity follows from $\hat{\rho}^{(0)}_{RI}$ being diagonal in the number-ket basis, and the anti-normally ordered characteristic function associated with the joint density operator for the $\{e^{i\phi}{\hat{\boldsymbol{a}}}_{R},{\hat{\boldsymbol{a}}}_{I}\}$ modes under hypotheses $H_{1}$ being given by Eq. (20) for all phase shifts $\phi$.

### v2

@@ Abstract
Quantum illumination (QI) is an entanglement-based protocol for improving lidar/radar detection of unresolved targets beyond what a classical lidar/radar of the same average transmitted energy can do. Originally proposed by Lloyd as a discrete-variable quantum lidar, it was soon shown that his proposal offered no quantum advantage over its best classical competitor. Continuous-variable, specifically Gaussian-state, QI has been shown to offer true quantum advantage, both in theory and in table-top experiments. Moreover, despite its considerable drawbacks, the microwave version of Gaussian-state QI continues to attract research attention. Recently, however, Pannu et al. (arXiv:2407.08005 [quant-ph]) have: (1) combined the entangled state from Lloyd’s QI with the channel models from Gaussian-state QI; (2) proposed a new positive operator-valued measurement for that composite setup; and (3) claimed that, unlike Gaussian-state QI, their QI achieves the Nair-Gu lower bound on QI target-detection error probability at all noise brightnesses. Pannu et al.’s analysis was asymptotic, i.e., it presumed infinite-dimensional entanglement. This paper works out the finite-dimensional performance of Pannu et al.’s QI. It shows that there is a threshold value for the entangled-state dimensionality below which there is no quantum advantage, and above which the Nair-Gu bound is approached asymptotically. Moreover, with both systems operating with error-probability exponents 1 dB lower than the Nair-Gu bound’s, Pannu et al.’s QI requires much higher entangled-state dimensionality than does Gaussian-state QI to achieve useful error probabilities in both high-brightness (100 photons/mode) and moderate-brightness (1 photon/mode) noise. Furthermore, neither system has appreciable quantum advantage in low-brightness ($\ll 1$ photon/mode) noise.
@@ I Introduction
Lloyd [5] coined the term “quantum illumination” for a lidar that transmitted a sequence of $M$-mode single-photon states while retaining their maximally-entangled single-photon companions for a joint measurement with the returned radiation. He assumed that the environment being probed never returned more than one photon in response to each transmission, and that returned photon was either background noise or a target reflection. Lloyd compared his QI performance with that of a single-photon (SP) lidar that probed the environment with the same state as his QI lidar, but had no stored idler. He argued that QI’s entanglement would make it difficult for a background photon to masquerade as the entangled companion of QI’s stored idler. Indeed, as compared to SP target detection, Lloyd’s QI target detection afforded a factor-of-$M$ improvement in error-probability exponent in his high-noise regime, i.e., when it is vastly more probable that the returned radiation from a single transmission is due to background as opposed to target reflection. That said, Lloyd’s QI and SP lidars are both quantum lidars, as they employ nonclassical transmitter states. Thus, when Shapiro and Lloyd [6] compared Lloyd’s QI with its best classical-lidar counterpart it turned out that the former could do no better than the latter, and could perform much worse. As a result, interest in Lloyd’s discrete-variable QI languished and was supplanted by interest in a continuous-variable version of QI, viz., Tan et al.’s Gaussian-state QI lidar [7].
@@
Pannu et al. do not assume that at most one photon is returned per transmission from the region being probed, hence avoiding the root cause of Lloyd’s QI not outperforming its best classical competitor. Their analysis is asymptotic, in that they pass to the limit $M\rightarrow\infty$ for the discrete-variable entangled state introduced in Lloyd’s QI. Doing so drives their receiver’s false-alarm probability to zero, and makes it simple to prove that their receiver realizes 6 dB quantum advantage in error-probability exponent in high-brightness noise, matching both Tan et al.’s performance in that regime and the Nair-Gu bound [13] on the attainable error-probability exponent of all possible QI protocols. Indeed, in this asymptotic regime, Pannu et al.’s QI matches the Nair-Gu bound at all noise brightnesses. However, contrary to their claim that Gaussian-state QI does not achieve Nair-Gu performance at low-noise brightness, it is easily shown [14] that the system’s error-probability exponent approaches the Nair-Gu bound, regardless of the noise brightness, as its signal brightness is decreased.
It is obvious that requiring infinite entangled-state dimensionality, i.e., $M\rightarrow\infty$, will put Pannu et al. QI beyond the realm of practicality, and Ref. [12] presents no finite-$M$ results for the error-probability exponent. Our paper remedies the preceding problem. In particular, we introduce a new, more explicit, result for the joint density operator of the returned and retained radiation when the target is present. Using this result we obtain accurate approximations to the single-shot false-alarm and detection probabilities for Pannu et al.’s receiver. From those approximations we then get the finite-$M$ multi-shot likelihood-ratio test and use the Chernoff bound to derive its error-probability exponent. We find that Pannu et al.’s QI has good and bad regimes, analogous to those of Lloyd’s QI, and that only in the good regime—which requires $M$ to exceed a threshold value—does it offer quantum advantage for the error-probability exponent.
@@
where the $\{\hat{a}_{B_{m}}:1\leq m\leq M\}$ are annihilation operators for background-noise modes that are in independent, identically-distributed (iid) thermal states with average photon number $N_{B}$ [16]. On the other hand, under hypothesis $H_{1}$ (target present), they use
@@
Pannu et al.’s QI thus offers a 6 dB quantum advantage in error-probability exponent when $N_{B}\gg 1$, no appreciable quantum advantage when $N_{B}\ll 1$, and 4.6 dB quantum advantage at $N_{B}=1$. Tan et al.’s QI matches those behaviors because, as noted earlier, its error-probability exponent approaches the Nair-Gu bound in the limit of low signal brightness [14]. See Sec. VI for a more detailed appraisal of Pannu et al. QI versus Tan et al. QI.
@@ V Multi-Shot Likelihood Ratio and Error-Probability Exponent
$\exp$$\{-[\kappa N_{T}/(N_{B}+1)]\mathcal{E}(\kappa,N_{B},M)\}$
@@
and focus our attention on two special cases, $\kappa=0.001$ with $N_{B}=100$ and $\kappa=0.001$ with $N_{B}=1$, as representatives of a weakly-reflecting embedded in either high-brightness or moderate-brightness noise [18].
It is easily verified that $\lim_{M\rightarrow\infty}\mathcal{E}(\kappa,N_{B},M)=1$, so that our error-probability exponent matches the Nair-Gu bound at infinite $M$, as shown by Pannu et al. But how high must $M$ be to approach that limit? Figure 1 plots $\mathcal{E}(\kappa,N_{B},M)$ versus $\log_{10}(M)$ for our representative cases, and Table 1 lists some key values therefrom. For the high-brightness noise we see that: (1) below the $M=7.34\times 10^{5}$ threshold, Pannu et al. QI offers no quantum advantage in error-probability exponent; and (2) $M=2.27\times 10^{13}$ is necessary for Pannu et al. QI’s error-probability exponent to be 1 dB lower than the Nair-Gu bound’s. Similarly, for the moderate-brightness noise we find that: (1) below the $M=7.33\times 10^{3}$ threshold, Pannu et al. QI offers no quantum advantage in error-probability exponent; and (2) $M=2.21\times 10^{11}$ is necessary for Pannu et al. QI’s error-probability exponent to be 1 dB lower than the Nair-Gu bound’s.
Figure 1: Plots of the penalty functions $\mathcal{E}(0.001,1,M)$ (top curve) and $\mathcal{E}(0.001,100,M)$ (bottom curve) versus $\log_{10}(M)$. Note that for both curves the $M$ values shown here exceed the penalty functions’ $M_{0}$ thresholds for $p_{D}>p_{F}$, viz., 22.6 for the moderate-brightness noise, and 31.5 for the high-brightness noise.
@@
$M$ $\mathcal{E}(0.001,1,M)$ Comment
$2.21\times 10^{11}$ $10^{-0.1}$ 1 dB off Nair-Gu bound
$7.33\times 10^{3}$ 0.25 no quantum advantage
Table 1: $M$ values needed for representative values of $\mathcal{E}(0.001,100,M)$ and $\mathcal{E}(0.001,1,M)$.
@@ VI Conclusions and Discussion
At this juncture, a comparison between finite-$M$ Pannu et al. QI and Tan et al. QI is warranted. Both systems can match the Nair-Gu bound on target-detection error-probability exponent for a weakly-reflecting ($\kappa\ll 1$) target embedded in thermal noise. Moreover, neither offers any appreciable quantum advantage for low-brightness ($N_{B}\ll 1$) background noise. That said, the conditions required for each of these protocols to realize their respective quantum advantages are quite different.
@@
Consider Pannu et al. QI and Tan et al. QI for our $\kappa=0.001$ with $N_{B}=100$ and $\kappa=0.001$ with $N_{B}=1$ examples, with both systems operating at error-probability exponents 1 dB lower than the Nair-Gu bound’s. In the high-brightness noise, Pannu et al.’s QI requires $M=2.27\times 10^{13}$ to operate at 1 dB below the Nair-Gu error-probability exponent, and achieves the Chernoff-bound performance shown in Fig. 2, whereas in the moderate-brightness noise $M=2.21\times 10^{11}$ suffices for those purposes. Tan et al.’s QI, on the other hand, requires the signal brightness to be $N_{S}=0.01523$ to operate at 1 dB below the Nair-Gu bound in the high-brightness noise, whereas $N_{S}=0.01421$ suffices for that purpose in the moderate-brightness noise. In both of those cases it achieves Fig. 2’s Chernoff-bound performance, where $N_{T}=MN_{S}$ is now the average number of transmitted signal photons.
Figure 3 compares the entanglement dimensionalities of the two QI systems for the parameters used in Fig. 2. Here we see that Pannu et al. QI requires more than $10^{5}$ times the dimensionality that suffices for Tan et al. QI in the high-brightness noise and more than $10^{4}$ times the dimensionality that Tan et al. QI requires in the moderate-brightness noise.
@@
Figure 2: Log-log plots of the Chernoff bound, $\Pr(e)$, versus $N_{T}$ for Pannu et al. QI and Tan et al. QI when $\kappa=0.001$ with $N_{B}=100$ (top curve) and $\kappa=0.001$ with $N_{B}=1$ (bottom curve), with both systems are operating at 1 dB below the Nair-Gu error-probability exponent. In the high-brightness noise, Pannu et al. QI requires $M=2.27\times 10^{13}$ to reach this operating point, whereas in the moderate-brightness noise, $M=2.21\times 10^{11}$ suffices for this purpose. In both brightnesses $N_{T}$ is the number of transmitted signal photons. For Tan et al. QI, this operating point requires $N_{S}=0.01523$ in the high-brightness noise and $N_{B}=0.01421$ in the moderate-brightness noise. In both of these cases $N_{T}$ is now the average number of transmitted signal photons.
@@
Figure 3: Log-log plots of $M_{\rm Pannu}/M_{\rm Tan}$ versus $\Pr(e)$ where $M_{\rm Pannu}$ and $M_{\rm Tan}$ are the entangled-state dimensionalities associated with the $\Pr(e)$ versus $N_{T}$ curves from Fig. 2.
@@
As if Pannu et al. QI’s requiring $10^{4}$-to-$10^{5}$ times the entangled-state dimensionality of Tan et al. QI to achieve the same error probability were not bad enough, it incurs an even larger disadvantage when we look at time-bandwidth product. In the scalar-wave, unresolved-target scenario that we are considering, only temporal degrees of freedom are available. Hence, a single pulse of dimensionality $M$ must have a time duration $T$ and bandwidth $W$ satisfying $M=TW$. Reference [3] explains that Tan et al.’s QI can carve $T=M/W$-s duration pulses from the signal and idler outputs of a continuous-wave-pumped parametric downconverter with $W$-Hz phase-matching bandwidth to obtain the $M$-dimensional entangled state it needs. It is not clear how to generate the $M$-dimensional entangled-state pulses that Pannu et al. QI needs, but a sequence of $N_{T}>10^{4}$ such pulses are needed for that QI system to realize the Chernoff-bound error probabilities $<10^{-3}$ from Fig. 2 in the moderate-brightness noise, and $N_{T}>10^{5}$ pulses are required in the high-brightness noise. Thus Pannu et al. QI’s total required time-bandwidth product is at least $10^{8}$-to-$10^{10}$ times what suffices for Tan et al.’s QI to get to $\Pr(e)<10^{-3}$ in our Fig. 2 examples.
@@
Our final points of comparison between these two QI protocols concern the four enormous hurdles, cited in Sec. I, that currently preclude finding a realistic target-detection use case for Tan et al.’s QI: its need to interrogate one resolution bin at a time; its need for a quantum memory to store its high time-bandwidth product idler; its need for extremely high time-bandwidth product radiation to get an acceptably low error probability; and its need for an interferometric measurement. How does Pannu et al.’s QI stack up against these hurdles? There is at least some encouragement in this regard, but so far the overall prospects are pretty poor. In particular, Pannu et al.’s QI still needs to interrogate one resolution bin at a time, and it still needs a quantum memory to store its high-dimensionality idler. However, assuming Pannu et al. QI’s single-pulse idler dimensionality is the same as that of Tan et al.’s QI, its discrete-variable memory may be easier to implement than Tan et al.’s continuous-variable memory. As we have seen, however, Pannu et al.’s required single-pulse idler dimensionality is apt to be $10^{4}$-to-$10^{5}$ times that of Tan et al.’s QI. Worse, Pannu et al. QI’s total time-bandwidth product may have to be a factor of $10^{8}$-to-$10^{10}$ times that of Tan et al. QI, owing to its need to transmit a long sequence of single-photon pulses. A final positive note for Pannu et al.’s QI is that it does not require an interferometric measurement. This phase insensitivity follows from $\hat{\rho}^{(0)}_{RI}$ being diagonal in the number-ket basis, and the anti-normally ordered characteristic function associated with the joint density operator for the $\{e^{i\phi}{\hat{\boldsymbol{a}}}_{R},{\hat{\boldsymbol{a}}}_{I}\}$ modes under hypotheses $H_{1}$ being given by Eq. (20) for all phase shifts $\phi$.